Assume that a person wants to download a one minute long sixty frames per second video . This frame is in 1920 times 1080 pixels with 3 bytes per pixel. Let us hypothesize the video is not caressed . How long does it take be downloaded this video over the channel ?!
Answer #1

Number of pixel = 1920*1080*3 = 6220800 bytes

i.e nearly 6MB

Frame rate = 6220800*60= 37324800 bytes/s ie 373 Mbs

Total time take= 373mbs/60 = 6.21 sec